OpenStudy (anonymous):

simlify it 4ae^2ax/e^4ax+2e^2ax+1 Should get this 4a/(e^ax+e^-ax)^2?

OpenStudy (anonymous):

e^4ax+2e^2ax+1=(e^2ax+1)^2 so deviding numerator and denominator by e^2ax you get the result

OpenStudy (anonymous):

yes yes i got it

OpenStudy (anonymous):

4a/e^-2ax(e^2ax+1)^2= 4a/[e^-ax(e^2ax+1)]^2=4a/(e^ax+e^-ax)^2